# LLMs for Targeted Sentiment in News Headlines: Exploring Different Levels of Prompt Prescriptiveness

### v1

@@ Abstract
News headlines often evoke sentiment by intentionally portraying entities in particular ways, making targeted sentiment analysis (TSA) of headlines a worthwhile but difficult task. Fine-tuned encoder models show satisfactory TSA performance, but their background knowledge is limited, and they require a labeled dataset. LLMs offer a potentially universal solution for TSA due to their broad linguistic and world knowledge along with in-context learning abilities, yet their performance is heavily influenced by prompt design. Drawing parallels with annotation paradigms for subjective tasks, we explore the influence of prompt design on the performance of LLMs for TSA of news headlines. We evaluate the predictive accuracy of state-of-the-art LLMs using prompts with different levels of prescriptiveness, ranging from plain zero-shot to elaborate few-shot prompts matching annotation guidelines. Recognizing the subjective nature of TSA, we evaluate the ability of LLMs to quantify predictive uncertainty via calibration error and correlation to human inter-annotator agreement. We find that, except for few-shot prompting, calibration and F1-score improve with increased prescriptiveness, but the optimal level depends on the model.
@@
Exploring Different Levels of Prompt Prescriptiveness
@@ 1 Introduction
News framing impacts information perception, shapes public opinion, and guides discussions on key topics Semetko and Valkenburg (2000). As succinct and attention-grabbing introductions to full news stories, headlines play a crucial role in news dissemination. They often evoke sentiment by portraying entities in specific ways, thus motivating the need for targeted sentiment analysis. Targeted sentiment analysis (TSA) is the task of determining the polarity of sentiment expressed towards the target entity Pei et al. (2019). While sentiment analysis inherently presents challenges due to its subjectivity, TSA introduces an additional layer of complexity by requiring the differentiation between targeted and overall sentiment. Analyzing the sentiment of news headlines is even more intricate because headlines are connotation-rich and culture-dependent, often involving implicit and indirect sentiment. This requires a detailed interpretive effort, including interpreting quotes, identifying negative events linked to entities, and navigating mixed-view headlines Hamborg et al. (2021).
@@
Fine-tuned encoder models such as BERT (Devlin et al., 2019) demonstrate strong TSA performance across various languages Wu and Ong (2021); Zhang et al. (2020); Mutlu and Özgür (2022). However, using these models in different languages or domains requires new fine-tuning, and adapting them to low-resource languages necessitates pre-trained models and labeled data. In contrast, large language models (LLMs) offer a reliable and versatile approach to TSA across various contexts by leveraging their broad linguistic and world knowledge, as well as in-context learning (Brown et al., 2020), without the need for annotated datasets or model fine-tuning. However, LLMs often lack the performance consistency of dataset-specific fine-tuned models. Moreover, LLMs’ performance is contingent on prompt design, posing a significant challenge in identifying optimal settings.
@@
In this paper, we begin by comparing the zero- and few-shot performance of open and closed-source LLMs to fine-tuned encoder models. We then explore the influence of prompt design on the performance of LLMs for TSA of news headlines, drawing parallels with annotation paradigms for subjective tasks. Similar to crafting effective annotation guidelines, successful prompt design for subjective tasks hinges on finding the appropriate level of prescriptiveness. Drawing from the descriptive-prescriptive dichotomy in linguistics and ethics, Rottger et al. (2022) proposed two contrasting paradigms for data annotation: the descriptive paradigm, which fosters annotator subjectivity and encourages diverse interpretations, and the prescriptive paradigm, which discourages subjectivity and mandates adherence to specific interpretation guidelines. The recent utilization of LLMs as data annotators (Wang et al., 2021; Pangakis et al., 2023; Alizadeh et al., 2023) invites a direct comparison between these paradigms and prompt design: a simple, less prescriptive prompt grants the LLM more freedom in interpreting the input, whereas a detailed, highly prescriptive prompt restricts the interpretation. Building on this parallel, we evaluate the predictive accuracy of state-of-the-art LLMs using prompts with different levels of prescriptiveness, ranging from plain zero-shot to elaborate few-shot prompts matching annotation guidelines.
@@
Another interesting parallel between annotation and prompting is label variation. Regardless of whether subjectivity is encouraged, some human label variation is inevitable in subjective tasks and may be leveraged for improving model performance Mostafazadeh Davani et al. (2022). Similarly, LLM inconsistency, typically viewed as a limitation, can diversify responses to emulate human label variation. Recent LLM uncertainty quantification methods Rivera et al. (2024); Xiong et al. (2023); Tian et al. (2023) can produce well-calibrated confidence scores, reflecting the task’s inherent subjectivity. Building on this, we assess LLMs’ capability to quantify predictive uncertainty in TSA of headlines using calibration error and correlation with human inter-annotator agreement.
@@
Our experiments use a Croatian dataset labeled with TSA on news headlines accompanied by detailed annotation guidelines. Additionally, we evaluate zero-shot LLMs and BERT on English and Polish TSA datasets. Our contributions include (1) comparing LLMs and BERT for TSA on Croatian, English, and Polish news headlines, (2) evaluating the effect of prompt prescriptiveness on LLMs’ predictive accuracy, and (3) assessing calibration error across models based on prompt prescriptiveness. This study offers valuable insights into LLMs’ zero- and few-shot potential for TSA of news headlines.
@@ 2 Related Work
Sentiment analysis on news headlines is an important task and garnered attention in prior work (Aslam et al., 2020; Agarwal et al., 2016; Rozado et al., 2022), especially for stock prediction (Joshi et al., 2016; Nemes and Kiss, 2021). In addition to overall sentiment, TSA, also known as fine-grained sentiment analysis, is crucial for understanding how entities are portrayed in news articles. Cortis et al. (2017) and Salgueiro et al. (2022) perform TSA on headlines from limited domains, namely financial and political headlines. Dufraisse et al. (2023) and Steinberger et al. (2011) present multilingual datasets for TSA in news articles. Hamborg and Donnay (2021) present a dataset for TSA on English news articles reporting on political topics, while (Balahur et al., 2013) focus on quotes from news articles. Overcoming the need for a labeled dataset, LLMs present as a possible solution for TSA due to their ICL abilities and broad background. Huang et al. (2020) conducted an analysis to identify and mitigate the entity bias of LLMs trained for sentiment analysis on Wikipedia and news articles. Chumakov et al. (2023) leverage both few-shot learning and fine-tuning with GPT models on mixed-domain Russian and English datasets to model sentiment effectively without domain-specific data.
@@ 3 Datasets and Models
Our experiments utilize the only two available datasets, to our knowledge, for TSA in general news headlines. These datasets cover three languages and employ different annotation styles.
@@
Filling the void in TSA for low-resource languages, the SToNe dataset Barić et al. (2023) offers overall sentiment and targeted sentiment annotations for Croatian news headlines, using ternary labels (positive, neutral, negative). Each of the 2855 headlines has six labels given by six annotators, and detailed annotator guidelines are provided. We leverage these guidelines to build prompts with increasing levels of prescriptiveness, with the highest level aligning with annotation guidelines and including provided examples.
@@
The SEN Baraniak and Sydow (2021) TSA dataset includes 3819 English and Polish news headlines. It comprises three parts: the Polish dataset (SEN_pl), the English dataset (SEN_en_r) annotated by volunteer researchers, and an English dataset (SEN_en_amt) annotated using Amazon Mechanical Turk. Unlike SToNe, SEN lacks raw labels, providing only an aggregated gold label for each headline (positive, neutral, and negative), and features less detailed and prescriptive annotator guidelines. Therefore, SEN is not suitable for our prompt prescriptiveness experiments.
In our experiments, we chose two open 7B models (Mistral and Neural Chat) derived from the LLama 2 model (Touvron et al., 2023), pitted against two proprietary OpenAI models (GPT-4 Turbo and GPT-3.5 Turbo) (OpenAI et al., 2023).
@@ 4 Experiments and Results
4.1 Accuracy Across Datasets
We first evaluate the LLMs’ accuracy of TSA on headlines and compare them to top-performing BERT models. We use the BERT models specifically pre-trained for each language – RoBERTa-base Liu et al. (2019) for English, BERTić Ljubešić and Lauc (2021) for Croatian, and Polish-RoBERTa-base-v2 Dadas et al. (2020) for Polish – and fine-tune each for TSA on the corresponding training set. For LLMs, we use zero-shot prompting on the test set, using basic prompts outlining the task and the target classes (cf. Appendix A).
@@ 4.1 Accuracy Across Datasets
STONE en_amt en_r pl
GPT 3.5 Turbo 61.3 66.1 61.5 60.0
GPT 4 Turbo 65.9 68.8 63.2 69.5
Mistral 43.0 56.1 45.8 47.7
Neural Chat 59.8 66.3 63.8 58.1
BERT* 77.9 63.6 56.2 61.9
@@
Table 1 presents the F1 scores on the test set portions of each dataset. GPT-4 Turbo achieves the highest F1 score on the Polish SEN and the crowdsourced English SEN. Interestingly, on the English SEN annotated by researchers, Neural Chat outperforms not only other fine-tuned BERT models but also GPT. On SToNe, BERTić surpasses all other models by a significant margin. This could be due to the highly prescriptive nature of the annotator guidelines, which dictate a specific interpretation of headline sentiment, a characteristic captured well by BERTić during fine-tuning. The relatively weaker performance of LLMs may stem from the prompts not aligning with that level of prescriptiveness – a question we explore next.
@@ 4.2 Level of Prompt Prescriptiveness
In this experiment, we utilize the SToNe dataset and its annotator guidelines to create six prompts of increasing prescriptiveness level. Each subsequent level incorporates additional information from the guidelines. Table 2 outlines these six levels (cf. Appendix A for full prompts). Our goal is to assess the ability of LLMs to follow instructions as accurately as human annotators and to determine the most effective level of prompt prescriptiveness.
@@
1
Concise, exploring the fundamental concepts of sentiment and targeted sentiment.
2
Includes a definition of targeted sentiment specifically within the framework of news headlines.
3
Provided with concise guidelines.
4
Comprehensive instructions provided as guidelines, excluding examples.
5
Comprehensive instructions presented as guidelines, including examples and brief explanations.
6
Comprehensive instructions provided exactly as they were presented to the annotators.
Table 2: Short descriptions of prompt prescriptiveness levels. Level 1 contains basic instructions, while level 6 adheres to the prescriptive paradigm and is identical to annotator guidelines (cf. Appendix A for full prompts).
Table 3 shows the results. We note a slight but notable variance in performance across all levels for all models. GPT-4 consistently outperforms other models across all levels, with GPT-4 and Neural Chat reaching their performance peaks at level 4 (detailed instructions formatted as guidelines without examples) and GPT 3.5 performing best at level 3 (concise guidelines). The performance drop seen at levels 5 and 6, the only ones with few-shot examples, may be due to the sensitivity regarding the selection and ordering of examples, a phenomenon observed in few-shot prompting (Lu et al., 2022; Chang and Jia, 2023). The trend of increased accuracy from levels 1 to 4 indicates a positive impact of providing LLMs with more prescriptive instructions. While fine-tuning BERT still outperforms prompting at these higher prescriptiveness levels, the attractiveness of using LLM for TSA on headlines lies in their universality across languages and the absence of a need for labeled datasets.
Level Neural Chat GPT 3.5 Turbo GPT 4 Turbo 1 59.8 60.1 65.9 2 61.2 58.3 64.3 3 61.5 65.7 69.9 4 63.1 64.0 70.2 5 60.5 63.0 66.8 6 62.5 64.5 68.2
@@ 4.3 Uncertainty Quantification
Given the inherent subjectivity of TSA and the stochastic nature of predictions generated by LLMs, we explore how the inherent uncertainty of LLM predictions can model human label variation and whether this varies across levels of prompt prescriptiveness. Using SToNe, we approach this question from two angles: (1) investigating if the uncertainty of LLM predictions aligns with inter-annotator disagreement and (2) examining the relationship between LLMs’ predictive and calibration accuracies.
@@
We use three uncertainty quantification methods: self-consistency sampling, distribution sampling, and verbal confidence assessment. Self-consistency sampling (SCS) (Xiong et al., 2023) leverages the inherent stochasticity of LLMs, influenced by internal parameters such as temperature. For each headline, we prompt the same model six times and accumulate the responses to mimic the distribution of six annotator responses, setting the temperature to 0.7 for all models. The second method, which we refer to as distribution sampling (DS), prompts the model to explicitly predict how six annotators would label the targeted sentiment, directly resulting in a distribution of positive, neutral, and negative responses. Lastly, the verbal confidence assessment (VCA) method (Xiong et al., 2023) prompts the LLM to produce three predictions for each headline, representing each sentiment class, along with a confidence score ranging from 0 to 100.
Figure 1 shows the correlations between label distribution entropies from LLMs and annotators across prompt prescriptiveness levels. While correlations vary, they are generally weak, indicating misalignment between LLM predictive uncertainty and human subjectivity. The highest correlation occurs with GPT 3.5 and DS at level 3.
Figure 1: Correlations between label distribution entropies from LLMs and human annotators, for various uncertainty quantification methods and across levels of prompt prescriptiveness
To analyze the calibration error, we consider only the labels with the highest confidence score for each headline and calculate the expected calibration error (ECE; cf. Appendix A.6). Figure 2 compares the LLMs predictive accuracy (F1 score) against calibration accuracy, defined as 1-ECE, with prompt prescriptiveness level indicated by dot size. Calibration accuracy is generally quite high (above $0.8$) across models and uncertainty methods. Similar to predictive accuracy, calibration accuracy tends to increase with prescriptivity level, except for the DS method of uncertainty quantification. Achieving a balance between F1 scores and calibration accuracy, the optimal configurations lie on the Pareto front. Among these, GPT 4 Turbo at prescriptiveness level 4 and Neural Chat at level 4 maximize the F1 score and calibration accuracy, respectively. GPT 3.5 Turbo at level 6 emerges as the best choice, achieving a balance between predictive and calibration accuracy.

### v3

# LLMs for Targeted Sentiment in News Headlines: Exploring the Descriptive–Prescriptive Dilemma

@@ Abstract
News headlines often evoke sentiment by intentionally portraying entities in particular ways, making targeted sentiment analysis (TSA) of headlines a worthwhile but difficult task. Due to its subjectivity, creating TSA datasets can involve various annotation paradigms, from descriptive to prescriptive, either encouraging or limiting subjectivity. LLMs are a good fit for TSA due to their broad linguistic and world knowledge and in-context learning abilities, yet their performance depends on prompt design. In this paper, we compare the accuracy of state-of-the-art LLMs and fine-tuned encoder models for TSA of news headlines using descriptive and prescriptive datasets across several languages. Exploring the descriptive–prescriptive continuum, we analyze how performance is affected by prompt prescriptiveness, ranging from plain zero-shot to elaborate few-shot prompts. Finally, we evaluate the ability of LLMs to quantify uncertainty via calibration error and comparison to human label variation. We find that LLMs outperform fine-tuned encoders on descriptive datasets, while calibration and F1-score generally improve with increased prescriptiveness, yet the optimal level varies.
@@
Exploring the Descriptive–Prescriptive Dilemma
@@ 1 Introduction
News headlines – succinct and attention-grabbing introductions to full news stories – often evoke sentiment by portraying entities in specific ways. Targeted sentiment analysis (TSA) is the task of determining the polarity of sentiment expressed towards the target entity Pei et al. (2019). While sentiment analysis is inherently challenging due to subjectivity, TSA introduces additional complexity by requiring the differentiation between targeted and overall sentiment.
For subjective tasks like TSA, the choice of data annotation paradigm is crucial. Rottger et al. (2022) identified two contrasting paradigms: descriptive and prescriptive. The descriptive paradigm encourages subjectivity and diverse interpretations, typically with brief guidelines. In contrast, the prescriptive paradigm discourages subjectivity by providing detailed interpretation guidelines.
@@
Fine-tuned encoders such as BERT (Devlin et al., 2019) show strong TSA performance across various languages Wu and Ong (2021); Zhang et al. (2020); Mutlu and Özgür (2022). However, using these models in different languages or domains requires new fine-tuning, and adapting them to low-resource languages necessitates pre-trained models and labeled data. In contrast, large language models (LLMs) offer a versatile approach to TSA across various domains by leveraging their broad linguistic and world knowledge, as well as in-context learning (Brown et al., 2020), without the need for annotated datasets or fine-tuning. However, LLMs performance is often inconsistent and contingent on prompt design (Mizrahi et al., 2024), making it challenging to identify optimal settings. Furthermore, it is unclear how specific TSA criteria, defined during annotation, can be transferred using zero- and few-shot prompting.
@@
In this paper, we compare the zero- and few-shot performance of open and closed-source LLMs to fine-tuned encoder models on datasets annotated following the descriptive or prescriptive paradigm. We then explore the influence of prompt design on the performance of LLMs for the prescriptive TSA dataset of Croatian news headlines. Similar to crafting effective annotation guidelines, finding the appropriate level of prescriptiveness is essential in prompt design. The recent use of LLMs as data annotators (Wang et al., 2021; Pangakis et al., 2023; Alizadeh et al., 2023) further invites a direct comparison of annotation paradigms and prompt design: less prescriptive prompts give more interpretive freedom, while highly detailed prompts constrain it. Building on this parallel, we evaluate the predictive accuracy of LLMs using prompts constructed from annotation guidelines with different levels of prescriptiveness, ranging from plain zero-shot to elaborate few-shot prompts matching annotation guidelines.
@@
Another interesting connection between annotation and prompting is label variation. Regardless of whether subjectivity is encouraged, some human label variation is inevitable in subjective tasks and may be leveraged to improve model performance Mostafazadeh Davani et al. (2022). Similarly, LLM inconsistency, typically viewed as a limitation, can diversify responses to emulate human label variation. Recent LLM uncertainty quantification methods Rivera et al. (2024); Xiong et al. (2023); Tian et al. (2023) can be used for the same purpose. Building on this idea, we assess LLMs’ capability to quantify predictive uncertainty in TSA of headlines using calibration error and compare label distribution with human label variation.
@@
Our experiments mainly focus on a Croatian dataset labeled with TSA on news headlines accompanied by detailed, prescriptive annotation guidelines. Additionally, we evaluate zero-shot LLMs and BERT on English, Polish, and Spanish TSA datasets with less prescriptive guidelines. Our contributions include (1) comparing LLMs and BERT for TSA on news headlines in four languages, (2) evaluating the effect of prompt prescriptiveness on LLMs’ predictive accuracy, and (3) assessing calibration error and label distribution across models based on prompt prescriptiveness. This study offers valuable insights into LLMs’ zero- and few-shot potential for TSA of news headlines.
@@ 2 Related Work
Sentiment analysis of news headlines is an important task that has garnered significant attention in prior work (Agarwal et al., 2016; Joshi et al., 2016; Aslam et al., 2020; Nemes and Kiss, 2021; Rozado et al., 2022). In addition to overall sentiment, TSA (aka fine-grained sentiment analysis) is crucial for understanding how entities are portrayed in news articles. Cortis et al. (2017) apply TSA on financial headlines, where sentiment is less implicit and topically constrained. Dufraisse et al. (2023) and Steinberger et al. (2011) present multilingual datasets for TSA in news articles. Hamborg and Donnay (2021) present a dataset for TSA on English news articles reporting on political topics, while (Balahur et al., 2013) focus on quotes from news articles. Overcoming the need for a labeled dataset, LLMs present as a possible solution for TSA due to their ICL abilities and broad background. Huang et al. (2020) conducted an analysis to identify and mitigate the entity bias of LLMs trained for sentiment analysis on Wikipedia and news articles. Chumakov et al. (2023) leverage both few-shot learning and fine-tuning with GPT models on mixed-domain Russian and English datasets to model sentiment effectively without domain-specific data.
@@ 3 Datasets and Models
Our experiments utilize, to our knowledge, the only two available datasets for TSA in general news headlines, alongside one domain-specific dataset. These datasets cover four languages and employ different annotation styles.
@@
The SToNe dataset Barić et al. (2023) offers overall sentiment and targeted sentiment annotations for Croatian news headlines, using ternary labels (positive, neutral, negative). Each of the 2855 headlines has 6 labels assigned by 6 annotators, with an IAA of $\kappa$ = 0.416 (moderate agreement). Annotators were instructed using prescriptive, detailed guidelines (obtained from the authors upon our request).
@@
The SEN Baraniak and Sydow (2021) dataset includes 3819 English and Polish news headlines. It comprises a Polish part (SEN_pl), an English part (SEN_en_r) annotated by volunteer researchers, and an English part (SEN_en_amt) annotated using Amazon Mechanical Turk. The reported Fleiss’ kappa IAA are $\kappa$ = .459, $\kappa$ = .309, and $\kappa$ = .303, respectively. Unlike SToNe, SEN lacks raw labels, providing only an aggregated gold label per headline (positive, neutral, and negative), and was annotated using vaguer annotator guidelines, adhering more to the descriptive paradigm.
Spanish.
[…]
Models.
We experiment with four open-source models: Neural Chat (NC) (7B), Llama 3 (8B), Phi-3 (3.8B), and Gemma (9B), pitted against two proprietary OpenAI models – GPT-4 Turbo (560B) and GPT-3.5 Turbo (175B) (OpenAI et al., 2023) (cf. Appendix A for more details).
@@ 4 Experiments and Results
4.1 Predictive Accuracy
We first evaluate the LLMs’ accuracy of TSA on headlines and compare them to top-performing BERT* models. We use the BERT models specifically pre-trained for each language – RoBERTa-base Liu et al. (2019) for English, BERTić Ljubešić and Lauc (2021) for Croatian, and Polish-RoBERTa-base-v2 Dadas et al. (2020) for Polish – and fine-tune each for TSA on the corresponding training set. For LLMs, we use zero-shot prompting on the test set, using basic prompts outlining the task and the target classes (cf. Appendix A).
@@ 4.1 Predictive Accuracy
STONE SD en_amt en_r pl
GPT 3.5 61.3 64.2 66.1 61.5 60.0
GPT 4 65.9 67.0 68.8 63.2 69.5
Neural Chat 59.8 63.0 66.3 63.8 58.1
Llama 3 53.5 60.5 59.2 52.7 51.2
Phi-3 43.5 61.7 58.3 52.7 47.3
Gemma 48.4 60.5 60.0 52.7 51.2
BERT* 74.9 66.7 63.6 56.2 61.9
@@
Table 1 presents the F1 scores on the test set portions of each dataset. On the descriptive datasets (SEN and Spanish), LLMs outperform BERT-based models. GPT-4 achieves the highest F1 score on the Polish SEN and the crowdsourced English SEN. Interestingly, on the English SEN annotated by researchers, NC outperforms both fine-tuned BERT models and GPT. However, on SToNe– the prescriptively annotated dataset – BERTić surpasses all other models by a significant margin. We argue this performance difference might stem from using different annotation paradigms. The best-performing LLMs seem to grasp the descriptive paradigm well, performing TSA closest to annotators. On the other hand, the performance gap observed in LLMs on SToNe may stem from the prompts’ vagueness and lack of alignment with its prescriptiveness – a question we explore next.
@@ 4.2 Level of Prompt Prescriptiveness
We utilize the SToNe dataset and its annotator guidelines to create six prompts of increasing prescriptiveness level, with each subsequent level incorporating additional information from the guidelines. Table 2 outlines these six levels (cf. Appendix A for full prompts). Our goal is to assess the ability of LLMs to follow instructions as accurately as human annotators and to determine the most effective level of prompt prescriptiveness.
@@
1 Concise, exploring the fundamental concepts of sentiment and targeted sentiment.
2 Includes a definition of targeted sentiment specifically within the framework of news headlines.
3 Provided with concise guidelines.
4 Comprehensive instructions provided as guidelines, excluding examples.
5 Comprehensive instructions presented as guidelines, including examples and brief explanations.
6 Comprehensive instructions provided exactly as they were presented to the annotators.
Table 2: Short descriptions of prompt prescriptiveness levels (cf. Appendix A for full prompts)
Table 3 shows the results. We observe variance in performance across all levels for all models. GPT-4 consistently outperforms other models across all levels, with GPT-4 and Neural Chat reaching their performance peaks at level 4 (detailed instructions formatted as guidelines without examples) and GPT 3.5 performing best at level 3 (concise guidelines). The performance drop seen at levels 5 and 6, the only ones with few-shot examples, may be due to the sensitivity regarding the selection and ordering of examples, a phenomenon observed in few-shot prompting (Lu et al., 2022; Chang and Jia, 2023). The increasing accuracy from levels 1 to 4 suggests that more prescriptive instructions positively impact LLM performance. Despite their overall lower performance, Llama 3, Gemma, and Phi-3 significantly improve at levels 5 and 6 (few-shot prompts). This difference in performance could be due to instruction tuning, which may have reduced sensitivity to few-shot configuration and improved context following.
Level NC GPT 3.5 GPT 4 Llama 3 Gemma Phi-3 1 59.8 60.1 65.9 53.5 48.4 43.5 2 61.2 58.3 64.3 50.9 48.8 40.6 3 61.5 65.7 69.9 52.9 55.8 44.2 4 63.1 64.0 70.2 51.9 53.1 43.6 5 60.5 63.0 66.8 60.6 59.3 49.4 6 62.5 64.5 68.2 61.9 59.4 46.3
@@ 4.3 Uncertainty Quantification
Given the inherent subjectivity of TSA and leveraging the stochastic nature of predictions generated by LLMs, we explore how LLMs can model human label variation and whether this varies across levels of prompt prescriptiveness. Using SToNe, we approach this question from two angles: (1) examining the relationship between LLMs’ predictive and calibration accuracies and (2) investigating if the uncertainty of LLM predictions aligns with inter-annotator disagreement.
@@
We use three uncertainty quantification methods: self-consistency sampling, distribution prompting, and verbal confidence assessment. Self-consistency sampling (SCS) (Xiong et al., 2023) leverages the inherent stochasticity of LLMs, influenced by internal parameters such as temperature. For each headline, we prompt the same model six times and accumulate the responses to mimic the distribution of six annotator responses, setting the temperature to 0.7 for all models. The second method, which we refer to as distribution prompting (DP), prompts the model to explicitly predict how six annotators would label the targeted sentiment, directly resulting in a distribution of positive, neutral, and negative responses. Lastly, the verbal confidence assessment (VCA) method (Xiong et al., 2023) prompts the LLM to produce three predictions for each headline, representing each sentiment class, along with a confidence score ranging from 0 to 100.
To analyze the calibration error, we consider only the labels with the highest confidence score for each headline and calculate the expected calibration error (ECE; cf. Appendix A.6). Figure 1 compares the LLMs predictive accuracy (F1 score) against calibration accuracy, defined as 1-ECE, with both metrics averaged across all models (cf. Tables 6 and 7 and Figure 3 in Appendix A.5 for comprehensive data across all models). As expected, as models’ predictions are evaluated solely against the gold labels, F1 is higher for SCS than for DP and VCA. Average calibration accuracy is generally high (above $0.75$) across models and uncertainty methods. Higher prescriptiveness levels show an increasing trend in both predictive and calibration accuracy, with the optimal level varying by uncertainty method (Level 6 for SCS and VCA, and Level 5 for DP). This suggests that using prescriptive annotation guidelines can enhance LLM performance for prescriptive datasets.
Figure 1: Average F1 scores and calibration accuracy for uncertainty quantification methods across levels (shaded ellipses indicate covariances)
Besides quantifying uncertainty, SCS and DP can model human label variation, implicitly (SCS) or explicitly (DP). We compare these label distributions to human label variation. Figure 2 shows heatmaps of average F1 scores for the two best-performing open- and closed-source models, GPT-4 and NC. The axes represent the majority vote per instance by annotators and model. The highest F1 score is achieved when both the annotators’ votes and the models’ prediction are unanimous (6 votes). The lowest F1 scores are generally achieved for instances with less agreement within annotators or model votes. For GPT-4, DP performs similarly to SCS, whereas for NC, there is a significant performance drop and dispersion of model votes in bins, signaling the model is not grasping the concept. This suggests that SCS is a better choice for modeling label distribution across models.
Figure 2: F1 score per majority vote bins for annotators (X) and model (Y) for SCS vs. DP for NC and GPT-4